OpenStudy (anonymous):

How many solutions does the equation 6n − 6n − 12 = 14 − 2 have? Two One None Infinite

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

I know it's not infinite but I think it's one

OpenStudy (holly00d1248):

wats 6n-6n

OpenStudy (anonymous):

n ?

OpenStudy (anonymous):

n gets cancelled out

OpenStudy (holly00d1248):

yes Ns cancelled and it makes it no solutions

OpenStudy (holly00d1248):

cancels

OpenStudy (anonymous):

oh

OpenStudy (igreen):

Yes.. 6n = 6n = 0 Anything subtracted from itself is 0.

OpenStudy (igreen):

That gives us: -12 = 14 - 2 Subtract: -12 = 12 -12 does not equal 12, so it has no solutions..if it was 12 = 12, it would be infinitely many solutions.

OpenStudy (igreen):

6n - 6n*

OpenStudy (anonymous):

Thank you guys

OpenStudy (holly00d1248):

Your welcome